You have one dollar bills and five dollar bills in your piggy bank. You know you have 12 bills all together. You take the money out and count the value of each bill. You discover that your total amount of money is 24 dollars. Write a system of equations. How many one dollar bills did you have? How many five dollar bills did you have?
Answer by Tomatoma(1)

# of 1 dollar bills = X
# of 5 dollar bills = Y
There is a total of 12 bills so:
X + Y = 12
The total amount of money is 24 dollars so:
Multiply X and Y by their $ values
1*X + 5*Y = 24
Change X + Y = 12 to X = 12 - Y
Substitute X in 1X + 5Y = 24
To get 12 - Y + 5*Y = 24
Then simplify that equation to get 4*Y = 12
Simplify further to get Y = 3
X + Y = 12
X + 3 = 12
X = 9
# of 1$ bills = 9
# of 5$ bills = 3